# FineFilter: A Fine-grained Noise Filtering Mechanism for Retrieval-Augmented Large Language Models

### v1

@@ 6 Conclusion
In this paper, we introduce FineFilter, a novel fine-grained noise filtering mechanism aimed at improving performance and reducing cost in RAG systems. By framing noise filtering as a sentence-level MinMax optimization problem, FineFilter effectively tackles the challenge of identifying relevant clues in complex reasoning scenarios. Its three optimized modules leverage KNN clustering to obtain sufficient relevant context and retain effective clues based on generator feedback. Experiments show that FineFilter outperforms baselines on three QA datasets in both performance and efficiency. Future work can explore more adaptive noise filtering that dynamically adjusts based on query complex or retrieval quality for complex reasoning tasks.
7 Limitations
@@
Appendix A Prompt
A.1 Prompt for Clue Extractor
A.2 Prompt for Adaptive Truncator
A.3 Prompt for Generator
Appendix B Details of Experimental Settings
@@
Appendix C Case Study
Prompt for Clue Extractor You are a highly skilled assistant specializing in extracting relevant information from provided documents. Your task is to identify and extract sentences from the documents as much as possible that are most directly useful for answering the given question. Rank the sentences in order of relevance, with the most relevant sentence listed first. Preface each sentence with its sequence number as follows: Sentence 1: …… Sentence n: Question: {Question} Documents: {Documents}
Table 6: Prompt for Clue Extractor.
Prompt for Adaptive Truncator You are a highly skilled assistant specializing in optimizing language model efficiency by truncating context based on question complexity and document quality. Given a question and a ranked list of sentences, identify and retain the most relevant ones while truncating the irrelevant sentences. Question: {Question}
Ranked List:
{Ranked List}
Table 7: Prompt for Adaptive Truncator.
Prompt for Generator [INST] <<SYS>>
@@ Appendix C Case Study
Table 8: Prompt for Generator.

### v4

@@ 7 Conclusion
We propose FineFilter, a fine-grained noise filtering mechanism to enhance performance and efficiency in RAG. By framing noise filtering as a sentence-level MinMax optimization problem, it effectively identifies relevant clues in complex reasoning scenarios. Its three optimized modules use KNN clustering to gather sufficient context and retain key clues based on generator feedback. Experiments show FineFilter outperforms baselines in performance and efficiency on three QA datasets. Future work can explore adaptive noise filtering that dynamically adjusts based on query complexity or retrieval quality for complex reasoning tasks.
Limitations
@@
Ethics Statement
This paper presents FineFilter, a fine-grained noise filtering mechanism that formulates RAG noise reduction as a sentence-level MinMax optimization problem. It extracts sufficient supporting clues, reranks them based on answerability, and adaptively truncates redundancy to enhance both answer accuracy and inference efficiency. Throughout this research, we have adhered to ethical guidelines to ensure the integrity and fairness of our work. All experiments are conducted using publicly available datasets, and the retrieval corpus is based on open-domain sources such as Wikipedia, ensuring transparency and reproducibility.
FineFilter does not involve any personally identifiable information or private user data. The proposed methods are designed to enhance model efficiency and robustness without reinforcing harmful biases. Nevertheless, as with all RAG systems, the quality and neutrality of retrieved content can influence outputs. Future work may incorporate bias detection and content verification modules to further improve fairness and reliability.
Appendix A More Details of Experimental Settings
@@
Appendix B Prompt
B.1 Prompt for the Generator
B.2 Prompt for the Clue Extractor
B.3 Prompt for the Adaptive Truncator
Prompt for the Generator [INST] <<SYS>>
@@ B.3 Prompt for the Adaptive Truncator
Table 8: Prompt for the Generator.
Prompt for the Clue Extractor You are a highly skilled assistant specializing in extracting relevant information from provided documents. Your task is to identify and extract sentences from the documents as much as possible that are most directly useful for answering the given question. Rank the sentences in order of relevance, with the most relevant sentence listed first. Preface each sentence with its sequence number as follows: Sentence 1: …… Sentence n: Question: {Question} Documents: {Documents}
Table 9: Prompt for the Clue Extractor.
Prompt for the Adaptive Truncator You are a highly skilled assistant specializing in optimizing language model efficiency by truncating context based on question complexity and document quality. Given a question and a ranked list of sentences, identify and retain the most relevant ones while truncating the irrelevant sentences. Question: {Question}
Ranked List:
{Ranked List}
Table 10: Prompt for the Adaptive Truncator.
Appendix C Additional Experimental Analysis
C.1 Impact of Different Rerankers
To select a more effective reranking base model, we compare BM25 Robertson et al. (2009), BGE-rerank Xiao et al. (2024), and Sentence-BERTReimers and Gurevych (2020). As shown in Table 11, the results demonstrate that Sentence-BERT outperforms all baseline models in both EM and F1 scores, so we chose it as the reranking base model.
Method EM F1
BM25 41.51 71.03
BGE-rerank 41.73 71.06
Sentence-BERT(Ours) 42.03 71.21
C.2 Weaker Extractor and Truncator
To better assess the impact of clue extractor and adaptive truncator quality on the end-to-end performance of RAG systems, we conduct an experiment in which FineFilter’s clue extractor and truncator are replaced with a weaker model, Flan-T5-Large. This setup simulates scenarios where the clue processing pipeline exhibits limited reasoning or compression capabilities, enabling us to examine the extent of performance degradation under such constraints.
As shown in Table 12, FineFilter consistently outperforms the baselines across all three datasets, although it still slightly underperforms compared to the Top-5 setting. This suggests that when the downstream generator is sufficiently powerful (e.g., LLaMA2-7B), relying solely on a simple or less capable clue processing module is insufficient to fully exploit the potential of the overall system. Instead, only upstream components with strong extraction and compression capabilities can interact effectively with a powerful generator to achieve high-quality question answering.
NQ TriviaQA HotpotQA
Method EM F1 EM F1 EM F1
LongLLMLingua 34.77 65.28 46.32 78.43 22.19 57.33
FILCO 31.96 64.05 37.71 73.24 19.76 55.74
RECOMP 35.12 68.71 42.96 77.39 23.91 58.19
BottleNeck 38.49 69.88 46.15 78.37 24.27 58.83
Top-1 36.81 69.21 42.74 77.13 25.54 60.09
Top-5 40.21 70.95 48.32 80.16 25.07 59.57
[…]
Table 12: Performance comparison on NQ, TriviaQA, and HotpotQA under a weakened setup where FineFilter’s clue extractor and adaptive truncator are replaced with Flan-T5-Large, to simulate limited clue processing capabilities.
NQ TriviaQA HotpotQA
Method EM F1 EM F1 EM F1
LongLLMLingua 35.71 68.89 46.28 78.31 22.06 57.13
FILCO 31.79 63.88 34.59 71.77 20.79 56.71
RECOMP 36.28 69.11 41.32 76.44 22.58 57.34
BottleNeck 37.94 69.83 45.91 77.94 23.26 58.02
Top-5 34.85 68.04 46.03 78.16 22.86 57.59
[…]
Table 13: Performance comparison on NQ, TriviaQA, and HotpotQA using Flan-T5-Large as the weaker generator, while keeping the clue processing pipeline unchanged.
C.3 Weaker Downstream Generator
To assess the impact of generator capacity on the overall performance of our RAG system, we replace the stronger generator (LLaMA2-7B) with a weaker model (Flan-T5-Large), while keeping all other components and configurations unchanged. We further conduct an ablation study on three datasets, as presented in Table 13 and Table 14.
Experimental results demonstrate that replacing a strong generator with a weaker model results in performance degradation across all datasets. Nevertheless, our method consistently outperforms the Top-5 baseline. Notably, the relative contribution of the adaptive truncator becomes more pronounced under weaker generation conditions Ding et al. (2025), particularly on reasoning-intensive datasets such as TriviaQA and HotpotQA. These findings indicate that, when the downstream generator is less capable, upstream components like the adaptive truncator play a more critical role. Therefore, in scenarios constrained by limited computational resources or smaller model sizes, well-designed filtering and compression strategies can effectively mitigate the limitations of weaker generators.
Method NQ TriviaQA HotpotQA
FineFilter 38.95 47.03 23.79
[…]
w/o clue reranker 36.91 45.33 22.94
w/o adaptive truncator 38.09 46.21 22.62
Table 14: Ablation study on NQ, TriviaQA, and HotpotQA test set using EM scores. The generator is Flan-T5-Large, and the clue processing pipeline is kept unchanged.
Method EM CR Latency (s.)
Ours (Truncator: LLaMA3-8B-Instruct) 42.17 $\boldsymbol{\times}$ 3.18
Ours (Truncator: Flan-T5-Large) 42.09 18.79$\times$ 3.05
Table 15: Performance of adaptive truncators with different parameter sizes on the NQ dataset.
C.4 Truncators with Different Parameters
To investigate the impact of truncator size on overall system performance, we replace the default truncator with a smaller Flan-T5-Large model and evaluate the system on the NQ dataset.
As shown in Table 15, using a smaller truncator leads to a slight drop in EM and CR, but provides a marginal improvement in inference speed (0.13s faster). These results suggest that smaller truncators can offer a reasonable trade-off between performance and efficiency. Therefore, we recommend choosing the truncator model based on application-specific requirements and resource constraints.
Appendix D Case Study